You and your friends are playing in the swimming pool with a 25-in diameter beach ball. How much force (in Newtons) would be needed to push the ball completely under water?

1
Expert's answer
2022-01-23T16:16:52-0500

Given:

"d=\rm 25\: in=0.635\: m"

"\rho=\rm 1000\: kg/m^3"

"g=\rm 9.8\:\rm N/kg"


The buoyant force

"F=g\rho V, \quad V=\frac{4}{3}\pi R^3"

Hence

"F=\rm 9.8*1000*4/3*3.14*(0.635/2)^3=1313\: N"